OpenStudy (anonymous):

What is the oxidation number of tin (Sn) in the compound Na2SnO2? -2 0 +2 +3

OpenStudy (abb0t):

sodium is just an ion which I don't think actually binds to tin, but is just there within close proximity to the molecule. So it might make sense if you sort of think of what might make this molecule neutral. meaning, zero. You know that sodium caries a +1 charge (but since there's 2, that means you have +2) Now, in order to make this equal to zero, what must the charge be of \(SnO_2\) be?

OpenStudy (anonymous):

oxidation number of sodium is 1 oxidation number of oxygen is -2 oxidation number of tin = x total oxidation number for compound is 0 hence, \[1 \times 2 + X + (-2) \times 2 = 0\] X = 4 - 2 = +2